Mrs. Y. is admitted to the Critical Care Unit with an acute myocardial infarction. Her physician...

  1. Mrs. Y. is admitted to the Critical Care Unit with an acute myocardial infarction. Her physician orders a Lidocaine drip via infusion pump. Order: Lidocaine 2 g in 500 ml D5W IV to infuse at 6 mg/min. What is the infusion rate in ml/min?
Answer #1

Given

Order:Lidocaine

Dose:2gm(2000mg)

Volume:500ml D5W

Rate:6mg/min

If 6mg= 1min

Then 2000mg=??min

Cross multiply and divide

=2000×1÷6

=2000÷6

=333.3 minutes (5.5 hours=333.3÷60)

Flow rate :Total volume (ml)×drop factor ÷Time×60

=500×15÷5.5×60

=500×15÷330

=7500÷330

=22.7 or 23 drops/ml

So to calculate ml/min

Divide 23÷ standard drop factor

=23÷15

=1.5 ml/min